OpenStudy (31356):

(b/7)^2 Show your work to apply the law of exponents! FREE MEDAL REWARDED!!

OpenStudy (31356):

Any ideas?

jimthompson5910 (jim_thompson5910):

Square each part that is in the parenthesis. Similar example (x/3)^2 = (x^2)/(3^2) (x/3)^2 = (x^2)/9

OpenStudy (31356):

Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

Tell me what you got so I can check

OpenStudy (31356):

Ok